OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 4x – 6 8x – 2y = 14 A. (14, 12) B. (12, 14) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

@PigPresident

OpenStudy (anonymous):

D no solution

OpenStudy (anonymous):

8x - 2(4x-6) = 14 8x-8x = 14 +12 no sol

OpenStudy (anonymous):

ok thanks